### v1

@@ Abstract
The Poisson probability distribution is frequently encountered in physical science measurements. In spite of the simplicity and familiarity of this distribution, there is considerable confusion among physicists concerning the description of results obtained via Poisson sampling. The goal of this paper is to mitigate this confusion by examining and comparing the properties of both conventional and popular alternative techniques. We concern ourselves in particular with the description of results, as opposed to interpretation. After considering performance with respect to several desirable properties we recommend summarizing the results of Poisson sampling with confidence intervals proposed by Garwood. We note that the $p$ values obtained from these intervals are well-behaved and intuitive, providing for a consistent treatment. We also find that averaging intervals can be problematic if the underlying Poisson distributions are not used. The considerations here provide a basis for describing results from other distributions as well.
@@
confidence set , descriptive statistics , Garwood interval , $p$ value , statistical coverage ,
@@ 1 Introduction
Poisson processes are familiar in physics. In searches for new phenomena, sampling from Poisson distributions with very few expected counts, including background, is common. We concern ourselves here with sampling random variable (RV) $N=0,1,2,\ldots$ from a Poisson distribution (pdf) of the form:
@@
where $n$ is a possible value of $N$, $\mu$ is the mean of the Poisson distribution, equal to the sum of the “signal strength” $\theta$, considered to be the unkown quantity of interest, and a background strength, $b\geq 0$, assumed here to be known. In fact it is often a function, $g(\theta)$, of the signal strength that we wish to learn about. We’ll assume for our discussion that $g(\theta)$ is strictly monotonic. This is the typical situation; other cases may require special treatment. It is also common that the background strength is only imperfectly known. However, it will suffice for our discussion to consider the case where $b$ is known, and the more general case can be considered in the context of our conclusions.
@@
We take it as our goal to describe the result of the measurement. Certainly, quoting the sampled value $n$, or equivalently nbn-b, is an excellent description of the measurement. It is a sufficient statistic for $\theta$ and we have an intuitive grasp of the Poisson RV. However, when we propagate this into a statement concerning a measurement of the physical quantity of interest, $g(\theta)$, this Poisson intuition tends to be obscured. This leads us to consider the more general framework of confidence intervals and $p$ values.
@@
Before embarking on this, it is necessary to consider more generally what we would like to accomplish. We have said that the goal is to describe the result of the measurement. This can be contrasted with what we ultimately would like to know: the value of $\theta$, or of $g(\theta)$. In order to avoid confusion, it is useful to distinguish statements about the measurement (sampled value $n$) from statements about the value of the parameter ($\theta$ or $g$), even though both statements may have expressions in terms of the parameter. We consider three aspects that we might be interested in: (i) point estimates; (ii) interval estimates in order to give an idea of how well determined is the point estimate; and (iii) hypothesis tests, for example to address whether a possible signal is significant.
@@
As mentioned above, we are ultimately interested in the value of the parameter. In this case, the point estimate can be regarded as what we believe to be the most probable value of the parameter. The word “believe” is important here – we don’t actually know the parameter value, so we can only say what we think (believe) is most likely. Likewise, an interval estimate assigns a probability with which we believe the parameter lies within a stated interval in parameter space. A hypothesis test assigns a probability to a statement about the value of the parameter. For example, we may have a 5% degree of belief that $\theta$ is greater than zero. It is important to understand that here we are making statements about the actual parameter value. This is the domain of Bayesian statistics.
@@
On the other hand, we here wish to describe the result of the measurement, without making any further inference about the value of the parameter. In this case, an appropriate point estimate could be simply $\hat{\theta}=n-b$, or $\hat{g}(\theta)=g(n-b)$. More generally, it could be computed as that value of the parameter for which the sampled value is most probable. A descriptive interval estimate could determine an interval in parameter space such that the interval contains the parameter value with a specified probabilty over the long run, that is, in the frequency interpretation of repeating the experiment many times. A hypothesis test has a similar interpretation. Here, are making no statement about the value of the parameter. For our Poisson sampling, we are only phrasing the fact that we observed $n$ events in different ways. If this appears to be trivial, it is – no matter how we say it, we are trying to say that we observed $n$ events. The utility of not just saying “nn” is in formulating the statement as a description connected with the physical parameter of interest. It also provides for a succinct standard framework for describing results independent of the sampling distribution.
@@
Given that what we care about is the actual parameter value, it seems reasonable to suggest that once we have made a measurement, there is no need to make descriptive statements if we can go ahead and quote the result in terms of statements about the paramter value. The problem is that we know of no way to accomplish this without doing so in the framework of degree of belief, and we have no right to expect that two people will agree to believe the same thing. Thus, we find merit in the idea of describing the result of the measurement in some objective way independent of consumer that is helpful for each to form their own inferences about the parameter value.
@@
There are many ways to describe something, including something as simple as the result of a measurement in a Poisson process. We have already mentioned giving the value of $n$, and g(nb)g(n-b). This is a good idea and should be done. However, we noted that this may not be so simple or complete a description when translated to the parameter of interest. We are typically interested in a point estimate, an estimate of variance, and if the Poisson process is a search for a signal of something new, an estimate of how significant our observation is compared with the no signal hypothesis. We’ll consider these in turn.
@@ 1.1 Point estimates
We may contrast this with the point estimator in a Bayesian (degree-of-belief) analysis in a situation where we know on physical grounds, for example, that $\theta$ cannot be negative. In this case we often (depending on the prior distribution) use the value of $\theta$ that maximizes the likelihood as our point estimate, but now we define the likelihood function to be zero for $\theta<0$, since we are certain that the (degree-of-belief) probability is zero for $\theta<0$. It may be supposed that we have incorporated the permitted range of $\theta$ into the prior distribution for $\theta$.
We regard our choice of point estimator for the Poisson model as obvious. Some might suggest that it not be permitted to be in a region where $\theta$ cannot lie. But this arises from a misunderstanding of the descriptive goal, perhaps as an attempt to make the estimate also serve as a statement about the true value. In particular, if we are aware that we have had a downward fluctuation in the background, it would not be a sensible description to obscure this fact. In general, it is a futile pursuit to try to simultaneously achieve description and interpretation.
@@ 1.2 Interval estimates
The point of an interval estimate in descriptive statistics is to provide information about how large the variations of the point estimate could be, that is, an idea of how far the point estimator may typically deviate from the expectation of the point estimator (the parameter value, assuming our point estimator is unbiased). A scale for such variation is the standard deviation, which in Poisson statistics is $\sqrt{\mu}$. Given a sampling $n$, an estimate for this is $\sqrt{n}$. This is indeed a fine choice. That is, we could quote an interval for $\theta$ of $(n-b-\sqrt{n},n-b+\sqrt{n})$, translating to an interval for $g(\theta)$ of $\left(g(n-b-\sqrt{n}),g(n-b+\sqrt{n})\right)$. The case when $n=0$ is not so much of a problem as one might suppose, since quoting an interval of zero size describes exactly the measurement.
@@
This problem is solved in the context of the frequency interpretation of statistics [3]: We make statements that are true with stated probability. That is, given any sampling $n$, we devise an algorithm to make a statement depending on $n$, such that the statement is true with the stated probability over (infinitely) repeated instances of the measurement. Sometimes people object that we generally don’t actually repeat the measurement, at least not a large number of times, and therefore this is unrealistic. But we don’t need to repeat the measurement to do the calculation, and the concept is well-defined in terms of the mathematics. This is the approach that is generally adopted for descriptive statistics and is the one that we shall develop further below. First, however, we’ll note other possible choices.
Another approach to a descriptive statistic may be derived from Bayesian methodology, even though it may thought of as such. We construct a posterior distribution for parameter $\theta$ from the likelihood function corresponding to observation $n$ according to:
@@
However, people often suggest prescribing such a prior based only on the sampling distribution. This is often called a reference or objective prior. This is an attempt to make a statement about the parameter without incorporating any prior knowledge or belief. Such an analysis proceeds as a Bayesian analysis in the sense that it may be interpreted as a statement of belief if one is willing to accept the reference prior as describing one’s prior degree of belief. There is no attempt, or desire, to satisfy a frequency interpretation. Biller and Oser, for example, have advocated this approach [4], based on uniform priors in some choice of parameterization. [^2] Some very nice discussion in this reference is confused, in my opinion, by interpreting frequency statistics as intended to serve a purpose beyond description. This is a common occurrence when referring to frequency statistics in discussions of Bayesian statistics.
@@
We may also consider “fiducial statistics” as a possible source of descriptive statistics. Fiducial statistics were invented to provide a framework for inference without the problem of the subjective prior of the Bayesian methodology [5]. The interpretation, and even the generality, of a “fiducial interval” is not clear; further discussion may be found, for example, in [6], [7], and [8]. Given certain conditions, a fiducial interval, $(\ell,u)$, corresponding to a given sampling $n$, is computed according to
$\begin{split}P(N\leq n|\theta=u(n))&=\alpha/2\\ P(N<n|\theta=\ell(n))&=1-\alpha/2.\end{split}$ (4)
Here, $1-\alpha$ is the “fiducial probability”, and we have equalized the probability in the two tails. We have also expressed this in notation in the context of our discussion of the Poisson distribution.
@@ 1.3 Signal significance
Besides describing a measurement in terms of a point estimate and an estimate of uncertainty, we are often interested in a description of how significant our observation is compared with the no signal hypothesis. As discussed above we’re going to adopt the frequency context to provide descriptive interval estimates with a probabilistic meaning. Thus, we’ll adopt the same context for statements of significance, again with a descriptive probabilistic meaning.
@@ 1.4 Following sections
It is crucial to understand that our discussion is in the context of frequency statistics. That is, we are only interested here in providing a description of a measurement. There is no attempt to make a statement about the true mean of the Poisson distribution. We suggest that we are providing a description that could be useful input towards forming such an inference, but we stop short of actually making an inference. We belabor this point because it is neglected in many discussions, contributing to ongoing confusion.
We keep our discussion focused on the basic question, without the complication of unknown nuisance parameters. The hope is that, once some basic issues are dealt with, then similar methodology may be applied in other situations. A discussion of this is timely, because of relatively recent advances in the understanding of some technical matters that we consider important.
We’ll continue in this introduction to develop the needed concepts. In section 2 we review the construction of frequency statistics and section 3 we consider potentially desirable properties. In sections 47 we discuss methods from the statistics community, followed by two methods invented in, and popular with, the particle physics community in section 8. Averaging of results is discussed in section 9, with discussion and conclusions in sections 10 and 11. We adopt notational conventions from ref. [9].
@@ 2 Construction of frequency statistics
To prepare for the following discussion, we briefly review the construction of frequency statistics. We begin with the notion of a confidence interval, or more generally a confidence set, followed by the construction of $p$ values, addressing significance.
@@ 2.1 Confidence sets
We are interested in confidence sets and $p$ values for a Poisson sampling process, for random variable (RV) $N=0,1,2,\ldots$ with distribution (pdf):
@@
A defining concept when discussing confidence sets is the “coverage”. That is, $C_{\alpha}(N)$ is a confidence set with confidence level $1-\alpha$ for parameter $\theta$ corresponding to RV $N$ iff
@@
The probability statement concerns random variable $C_{\alpha}(N)$ (independent of $\theta$), and we are working with the frequentist interpretation. This statement must hold for any possible value of $\theta$ (which need not be the same for repeated samplings). The quantity $1-\alpha$ is referred to as the “confidence level” of our set. If there is equality in Eq. 6, then the coverage of our confidence set is precise. [^3] This is also referred to as “exact” coverage, but we choose the word “precise” in order to avoid confusion with the use of “exact” introduced in section 3.1. Otherwise it “overcovers”. For discrete distributions such as the Poisson, obtaining precise coverage can be achieved by introducing an additional continuous RV [10, 11, 12], under the notion of a “randomized confidence set”. However, this does not add anything of essential value to our description, and we don’t advocate this. Thus, for the Poisson, we don’t insist on equality.
@@
That is, we construct a “confidence interval”.
2.2 $p$ values
In addition to describing the result of a measurement with a confidence interval, it can be useful to quote a $p$ value, for example, when we are searching for evidence of something “new”. The $p$ value provides a probability statement addressing how consistent a measurement is with a given hypothesis, in the context of a set of possible alternative hypotheses. Thus, we consider the question of consistency with a given value of the Poisson mean:

### v2

@@ Abstract
The Poisson probability distribution is frequently encountered in physical science measurements. In spite of the simplicity and familiarity of this distribution, there is considerable confusion among physicists concerning the description of results obtained via Poisson sampling. The goal of this paper is to mitigate this confusion by examining and comparing the properties of both conventional and popular alternative techniques. We concern ourselves in particular with the description of results, as opposed to interpretation. After considering performance with respect to several desirable properties we recommend summarizing the results of Poisson sampling with confidence intervals proposed by Garwood. We note that the $p$-values obtained from these intervals are well-behaved and intuitive, providing for consistent treatment. We also find that averaging intervals can be problematic if the underlying Poisson distributions are not used.
@@
confidence set , descriptive statistics , Garwood interval , $p$-value , statistical coverage ,
@@ 1 Introduction
Poisson processes are familiar in physics. In searches for new phenomena, sampling from Poisson distributions with very few expected counts, including background, is common. We concern ourselves here with sampling [^2] I consider “sampling” from a probability distribution to be synonymous with making an “observation”. random variable (RV) $N=0,1,2,\ldots$ from a Poisson distribution (pdf) of the form:
@@
where $n$ is a possible value of $N$, $\mu$ is the mean of the Poisson distribution, equal to the sum of the “signal strength” $\theta$, considered to be the unknown quantity of interest, and a background strength, $b\geq 0$, assumed here to be known. In fact it is often a function, $g(\theta)$, of the signal strength that we wish to learn about. We will assume for our discussion that $g(\theta)$ is strictly monotonic. This is the typical situation; other cases may require special treatment. It is also common that the background strength is only imperfectly known, or that there are other uncertainties. However, it will suffice for our discussion to consider the case where $b$ is known. We deliberately restrict to a simple case because there is already much to understand and even here there is confusion.
@@
We take it as our goal to describe the result of the measurement (we identify the term “measurement” as a sampling from a probability distribution). Certainly, quoting the sampled value $n$, or equivalently nbn-b, is an excellent description of the measurement. It is a sufficient statistic for $\theta$ and we have an intuitive grasp of the Poisson RV. However, when we propagate this into a statement concerning a measurement of the physical quantity of interest, $g(\theta)$, this Poisson intuition tends to be obscured. This leads us to consider the more general framework of confidence intervals and $p$-values.
@@
Before embarking on this, it is necessary to consider more generally what we would like to accomplish. We have said that the goal is to describe the result of the measurement. This can be contrasted with what we ultimately would like to know: the value of $\theta$, or of $g(\theta)$. In order to avoid confusion, it is useful to distinguish statements about the measurement (sampled value $n$) from statements about the value of the parameter ($\theta$ or $g$), even though both statements may have expressions in terms of the parameter. We consider three aspects that we might be interested in: (i) point estimates; (ii) interval estimates in order to give an idea of how well determined is the point estimate; and (iii) hypothesis tests, for example to address whether a possible signal could be significant.
@@
As mentioned, we are ultimately interested in the value of the parameter. In this case, the point estimate might be regarded as what we believe to be the most probable value of the parameter. The word “believe” is important here – we do not actually know the parameter value, so we can only say what we think (believe) is most likely. Likewise, an interval estimate assigns a probability with which we believe the parameter lies within a stated interval in parameter space. A hypothesis test assigns a probability to a statement about the value of the parameter. For example, we may have a 5% degree of belief that $\theta$ is greater than zero. It is important to understand that here we are making statements about the actual parameter value. This is the domain of Bayesian statistics.
@@
On the other hand, we here wish to describe the result of the measurement, without making any further inference about the value of the parameter. In this case, an appropriate point estimate could be simply $\hat{\theta}=n-b$, or $\hat{g}(\theta)=g(n-b)$. More generally, it could be computed as that value of the parameter for which the sampled value is most probable. A descriptive interval estimate could determine an interval in parameter space such that the interval contains the parameter value with a specified probability over the long run, that is, in the frequency interpretation of repeating the experiment many times. A hypothesis test has a similar interpretation. Here, we are making no statement about the value of the parameter. For our Poisson sampling, we are only phrasing the fact that we observed $n$ events in different ways. If this appears to be trivial, it is – no matter how we say it, we are trying to say that we observed $n$ events. The utility of not just saying “nn” is in formulating the statement as a description connected with the physical parameter of interest. It also provides for a succinct standard framework for describing results independent of the sampling distribution.
@@
Given that what we care about is the actual parameter value, it seems reasonable to suggest that once we have made a measurement, there is no need to make descriptive statements if we can go ahead and quote the result in terms of statements about the parameter value. The problem is that we know of no way to accomplish this without doing so in the framework of degree of belief, and we have no guarantee that two people will agree to believe the same thing. Thus, we find merit in the idea of describing the result of the measurement in some objective way, independent of consumer, which is helpful for each to form their own inferences about the parameter value.
@@
There are many ways to describe something, including something as simple as the result of a measurement in a Poisson process. We have already mentioned giving the value of $n$, and g(nb)g(n-b). This is a good idea and should be done. However, we noted that this may not be so simple or complete a description when translated to the parameter of interest. We are typically interested in a point estimate, an estimate of variance, and if the Poisson process is a search for a signal of something new, an estimate of how significant our observation is compared with the no signal hypothesis. We consider these in turn.
@@ 1.1 Point estimates
We may contrast this with the point estimator in a Bayesian (degree-of-belief) analysis in a situation where we might know on physical grounds, for example, that $\theta$ cannot be negative. In this case we often (depending on the prior distribution) use the value of $\theta$ that maximizes the likelihood as our point estimate, but now we define the likelihood function to be zero for $\theta<0$, since we are certain that the (degree-of-belief) probability is zero for $\theta<0$. It may be supposed that we have incorporated the permitted range of $\theta$ into the prior distribution for $\theta$.
We regard our choice of point estimator for the Poisson model as obvious. Some might suggest that it not be permitted to be in a region where $\theta$ cannot lie. This can arise from a misunderstanding of the descriptive goal, perhaps as an attempt to make the estimate also serve as a statement about the true value. In particular, if we are aware that we have had a downward fluctuation in the background, it would not be a sensible description to obscure this fact. In general, it is a futile pursuit to try to simultaneously achieve description and interpretation. See section 3 for further discussion.
@@ 1.2 Interval estimates
The point of an interval estimate in descriptive statistics is to provide information about how large the variations of the point estimate could be, that is, an idea of how far the point estimator may typically deviate from the expectation of the point estimator (the parameter value, assuming our point estimator is unbiased). A scale for such variation is the standard deviation, which in Poisson statistics is $\sqrt{\mu}$. Given a sampling $n$, an estimate for this is $\sqrt{n}$. This is indeed a fine choice. That is, we could quote an interval for $\theta$ of $n-b-\sqrt{n},n-b+\sqrt{n}]$, translating to an interval for $g(\theta)$ of $\left[g(n-b-\sqrt{n}),g(n-b+\sqrt{n})\right]$. The case when $n=0$ is not so much of a problem as one might suppose, since quoting an interval of zero size describes exactly the measurement.
@@
This problem is solved in the context of the frequency interpretation of statistics [27]: We make statements that are true with stated probability. That is, given any sampling $n$, we devise an algorithm to make a statement depending on $n$, such that the statement is true with the stated probability over (infinitely) repeated instances of the measurement. Sometimes people object that we generally do not actually repeat the measurement procedure, at least not a large number of times, and therefore this is unrealistic. But we do not need to repeat the measurement to do the calculation, and the concept is well-defined in terms of the mathematics. This is the approach that is generally adopted for descriptive statistics and is the one that we shall develop further below. First, however, we will note other possible choices.
Another approach to a descriptive statistic may be derived from Bayesian methodology, even though it may not be thought of as such. We construct a posterior distribution for parameter $\theta$ from the likelihood function corresponding to observation $n$ according to:
@@
However, people often suggest prescribing such a prior based only on the sampling distribution. This is often called a reference or objective prior. This is an attempt to make a statement about the parameter without incorporating any prior knowledge or belief. Such an analysis proceeds as a Bayesian analysis in the sense that it may be interpreted as a statement of belief if one is willing to accept the reference prior as describing one’s prior degree of belief. There is no attempt, or desire, to satisfy a frequency interpretation. Biller and Oser, for example, have advocated this approach [3], based on uniform priors in some choice of parameterization.
@@
We may also consider “fiducial statistics” as a possible source of descriptive statistics. Fiducial statistics were invented to provide a framework for inference without the problem of the subjective prior of the Bayesian methodology [14]. The interpretation, and even the generality, of a “fiducial interval” is not clear; further discussion may be found, for example, in [29], [49], and [21]. Given certain conditions, a fiducial interval, $\ell,u]$, corresponding to a given sampling $n$ of random variable $N$, is computed according to
$\begin{split}\hbox{Prob}[N\leq n|\theta=u(n)]&=\alpha/2\\ \hbox{Prob}[N\geq n|\theta=\ell(n)]&=\alpha/2.\end{split}$ (4)
Here, $1-\alpha$ is the “fiducial probability”, and we have equalized the probability in the two tails. We have here expressed this in notation in the context of our discussion of the Poisson distribution.
@@ 1.3 Signal significance
Besides describing a measurement in terms of a point estimate and an estimate of uncertainty, we are often interested in a description of how significant our observation is compared with the no signal hypothesis or alternatively, how consistent the data is with the no signal hypothesis. As discussed above we’re going to adopt the frequency context to provide descriptive interval estimates with a probabilistic meaning. Thus, we will adopt the same context for statements of significance, again with a descriptive probabilistic meaning.
@@ 1.4 Following sections
It is crucial to understand that our discussion is in the context of frequency statistics. That is, we are only interested here in providing a description of a measurement. There is no attempt to make a statement about the true mean of the Poisson distribution. We suggest that we are providing a description that could be useful input towards forming such an inference, but we stop short of actually making an inference. We belabor this point because it is neglected in many discussions, contributing to considerable confusion.
We keep our discussion focused on the basic question, without the complication of unknown nuisance parameters. The hope is that, once some basic issues are dealt with, then similar methodology may be applied in other situations. A discussion of this is timely, because of relatively recent advances in the understanding of some technical matters that we consider important, as will be discussed below.
We will continue to develop the needed concepts. In section 2 we review the construction of frequency statistics. Some controversial points concerning frequency statistics are discussed in section 3. In section 4 we consider potentially desirable properties. In sections 58 we discuss methods from the statistics community, followed by two methods invented in, and popular with, the particle physics community in section 9. Averaging results is discussed in section 10, with discussion and conclusions in sections 11 and 12.
We adopt notational conventions from ref. [25]. Throughout, $b$ is the known expected “background” counts and $\theta$ is the unknown expected signal counts, with $\mu\equiv\theta+b$. Corresponding hypothetical values of $\theta$ and $\mu$ are denoted by $\theta_{0}$ and $\mu_{0}$, respectively (and a “1” subscript denotes an alternative hypothetical value). In general, we nest our parentheses, $\{[()]\}$. We use $E()$ to denote an expectation value. We will generally denote our confidence intervals as closed intervals, following the convention in [38], though this is not universal (e.g., [1]), and it does not usually matter when dealing with a continuous parameter space. Exceptions will have $($ or )) where they occur. Sets are depicted with braces $\{\}$.
@@ 2 Construction of frequency statistics
To prepare for the following discussion, we briefly review the construction of frequency statistics. We begin with the notion of a confidence interval, or more generally a confidence set, followed by the construction of $p$-values, addressing significance.
@@ 2.1 Confidence sets
We are interested in confidence sets and $p$-values for a Poisson sampling process, for random variable (RV) $N=0,1,2,\ldots$ with distribution (pdf):
@@
A defining concept when discussing confidence sets is the “coverage”. That is, $C_{\alpha}(N)$ is a confidence set with confidence level (CL) $1-\alpha$ for parameter $\theta$ corresponding to RV $N$ iff
@@
The probability statement concerns random variable $C_{\alpha}(N)$ (independent of $\theta$), and we are working with the frequentist interpretation. This statement must hold for any possible value of $\theta$ (which need not be the same for repeated samplings). If there is equality in Eq. 6, then the coverage of our confidence set is precise. [^3] This is also referred to as “exact” coverage, but we choose the word “precise” in order to avoid confusion with the use of “exact” introduced in section 4.1. Otherwise, it “overcovers”. For discrete distributions such as the Poisson, obtaining precise coverage can be achieved by introducing an additional continuous RV [42, 9, 32], under the notion of a “randomized confidence set”. However, this does not add anything of essential value to our description, and we do not advocate this. Thus, for the Poisson, we do not insist on equality.
@@
That is, we construct a “confidence interval” at confidence level $1-\alpha$.
2.2 $p$-values
In addition to describing the result of a measurement with a confidence interval, it can be useful to quote a $p$-value ([38, 50]), for example, when we are searching for evidence of something “new”. The $p$-value provides a probability statement addressing how consistent a measurement is with a given hypothesis, in the context of a set of possible alternative hypotheses. Thus, we consider the question of consistency with a given value of the Poisson mean: